OpenStudy (anonymous):

which pair of slopes represent perpendicular lines? a. 1/2, -1/2 b. 1/3, 3 c. -1/4, 4 d. -1/5,-5 Best answer and follow :) thank u

rishavraj (rishavraj):

if product of slopes = -1 ...it means lines r perpendicular

OpenStudy (anonymous):

how would i get the product?

rishavraj (rishavraj):

like if m1 and m2 are slopes of two lines then if \[m_1 \times m_2 = -1\]

OpenStudy (anonymous):

im lost /:

rishavraj (rishavraj):

like see \[\frac{ 1 }{ 2 } \times -\frac{ 1 }{ 2 } = -\frac{ 1 }{ 4 }\]

rishavraj (rishavraj):

so its not perpendicular...bcoz product isn't equal to -1

rishavraj (rishavraj):

likewise check other options..... if the product is -1 ...then the lines are perpendicular

OpenStudy (anonymous):

how would i do 1/3 x 3?

rishavraj (rishavraj):

\[\frac{ 1 }{ 3 } \times 3 = 1\]

OpenStudy (anonymous):

so -1/4 x 4 would be -1?

rishavraj (rishavraj):

yup :))

rishavraj (rishavraj):

which means tht option is ur answer:))

OpenStudy (anonymous):

yes thank you so much once again for helping!!

rishavraj (rishavraj):

yw :))